Solve the following problems by USING Lagrange multipliers.

(a) Find the maximum and minimum values of f(x, y, z) = x^2 + y^2 + z^2 subject to the constraint (x − 1)^2 + (y − 2)^2 + (z − 3)^2 = 4

(b) Find the maximum and minimum values of f(x, y, z) = x^2 + y^2 + z^2 subject to the constraints (x − 1)^2 + (y − 2)^2 + (z − 3)^2 = 9 and x − 2z = 0.
Answer #1


(a) F(x.y,z)x2 +z A- +-2-3-4) F 0 2x +2A (x-1)0 _ (-1) F 0 = 2y+22 (y-2)0 _ (y-2) F 0 _ 2z 22(z-3) 0 -2 3 7- y Z = k z-3 (x-1

F (x.y.z) +) +(x-1(y2 4(s-3 - 2 +(x-2z) F0 2х +21(х-1)+ д%3D0; F0 2y+2A(y-2) 0, F 0 2z+2A (z-3)-24 0; 4x +4 (x-)+2u = 0; 2z+2

THUMB UP PLEASE